Yolanda invested a portion of $14,000 at 7.25% interest and the balance at 9.5% interest. How much did Yolanda invest at each rate if her total income from both investments was $1240 after one year?
Answer by tutorcecilia(2105)


7.25%x+9.5($14,000-x)=$1240
.0725x+.095(14,000-x)=1240
.0725x+1330-.095x=1240
.0725x-.095x=-1330+1240
-.0225x=-90
x=$4,000 invested at 7.25%
($14,000-4,000)=$10,000 invested at 9.5%
.
Check by plugging all of the values back into the original equation and solve.